OpenStudy (anonymous):

Find the answer using significant figures: [(2.46×10^6)÷45.637]+130.56

OpenStudy (itsbribro):

PEMDAS so parentheses and exponents first so what is \[10^6\]

OpenStudy (itsbribro):

10^6= 1,000,000

OpenStudy (itsbribro):

so now we have parentheses so:\[[(2.46\times1,000,000)\div45.637]+130.56\]

OpenStudy (itsbribro):

2.46 * 1,000,000 = 2460000

OpenStudy (itsbribro):

so now you have:\[[2,460,000\div45.637]+130.56\]

OpenStudy (itsbribro):

what is 2,460,000 divided by 45.637

OpenStudy (itsbribro):

it equals 53903.6308

OpenStudy (itsbribro):

not 53903.6308+130.56= 54034.1908

OpenStudy (itsbribro):

*im sorry not is supposed to be now

OpenStudy (itsbribro):

so i believe 54034.1908 is your answer. I'm sorry if this didn't help but if it did can i have a medal? thanks ^.^ byeee